# Weighted algebraic topology, II

### 05 May 2026 v1

@@ 2.5. Lipschitz maps
We say that a mapping $f\colon X\to Y$ between $\mu$-spaces is $\lambda$-Lipschitz, or a $\lambda$-map, if
(33) $\lambda\mu_{X}(x,x^{\prime})\geqslant\mu_{Y}(f(x),f(x^{\prime}))\qquad\qquad(x,x^{\prime}\in X),$
and we say that $f$ is a Lipschitz map, or $\mu_{\infty}$-map, if it is has some Lipschitz constant $\lambda\geqslant 0$, in the previous sense. Composing a $\lambda$-map with a consecutive $\lambda^{\prime}$-map we get a $\lambda\lambda^{\prime}$-map.
@@
The category $\mu\mathsf{Mtr}$ is a wide subcategory of the category $\mu_{\infty}\mathsf{Mtr}$ of $\mu$-spaces and Lipschitz maps. A Lipschitz isomorphism will be an isomorphism of the latter. The symmetric case is denoted as $\mu_{\infty}\mathsf{Mtr}$.
@@
(b) If $X$ is positive (a $\delta$-space) and $Y$ an arbitrary $\mu$-space, one can define, for every mapping $f\colon X\to Y$ (of sets) a Lipschitz weight, or seminorm $f||\in[0,\infty]$
(34) $||f||={\rm inf}\{\lambda\in[0,\infty[\,|\,\forall\;x,x^{\prime}\in X,\;\mu_{Y}(fx,fx^{\prime})\leqslant\lambda\delta_{X}(x,x^{\prime})\},$
@@
The categories $\delta_{\infty}\mathsf{Mtr}$ and $\delta_{\infty}\mathsf{Mtr}$ are multiplicatively weighted by the Lipschitz weight. By this we mean that every morphism $f$ is assigned a weight $||f||\in[0,\infty[$ so that:
@@
(c) The category $\mu_{\infty}\mathsf{Mtr}$ (resp. $\mu_{\infty}\mathsf{Mtr}$, $\delta_{\infty}\mathsf{Mtr}$, $\delta_{\infty}\mathsf{Mtr}$) is finitely complete and cocomplete, and the inclusion of $\mu\mathsf{Mtr}$ (resp. $\mu\mathsf{Mtr}$, $\delta\mathsf{Mtr}$, $\delta\mathsf{Mtr}$) preserves finite limits and colimits. But the ‘metric’ of a (co)limit in the larger category is only determined up to Lipschitz isomorphism.
If $X$ is a $\mu$-metric space and $\lambda\in[0,\infty[$, we shall write $\lambda X$ the same set equipped with the $\mu$-metric $\lambda\mu_{X}$. Thus, a $\lambda$-Lipschitz map is the same as a $\mu$-map $\lambda X\to Y$. This transformation preserves symmetric structures and positive structures.
@@ 2.6. Remarks
The extension to relative metrics brings about aspects which strongly diverge from the positive case. We begin with some elementary observations. In particular, they show that the structure of $\mu\mathbb{R}$ strongly restrains the Lipschitz maps defined on this space; in fact, we are more interested in maps from $\delta$-spaces to $\mu\mathbb{R}$ (as in 4.6).
(a) A $\mu_{\infty}$-map $f\colon\mu\mathbb{R}\to\mu\mathbb{R}$ is simply an affine map $f(x)=\lambda x+y_{0}$ (with $\lambda\geqslant 0$), and admits a unique Lipschitz constant, namely $\lambda$. (This will be extended to linear $\mu$-spaces, in 3.2(d).)
@@
(b) Any $\lambda$-Lipschitz maps $f\colon\mu\mathbb{R}\to\delta_{0}\mathbb{R}$ is constant, and $\lambda$ is necessarily 0 (see 3.2(e)).
@@
(c) Points as maps. A point of a $\mu$-space $X$ can be identified with a $\mu$-map $\delta\top\to X$ defined on the $\delta$-singleton. Formally, the forgetful functors $\mu_{\infty}\mathsf{Mtr}\to\mathsf{Set}$ and $\mu\mathsf{Mtr}\to\mathsf{Set}$ are represented by the $\delta$-singleton (the identity of the tensor product, see 6.1). If $X$ is positive every constant mapping $f\colon X\to Y$ is a $\mu$-map: it can be factorised as $X\to\delta\top\to Y$.
On the other hand, a $\mu_{\infty}$-map $\mu\top\to Y$ can only reach a flat point of $Y$. (The $\mu$-singleton represents the ‘flat-part functor’ mentioned in 2.2(b).)
@@ 2.7. Future and past topologies
A $\mu$-space $X=(X,\mu)$ can be equipped with several topologies, which coincide in the symmetric case.
@@
(36) $B^{+}(x_{0},\varepsilon)=\{x\in X\,|\,\mu(x_{0},x)<\varepsilon\}\qquad\;\;(\varepsilon>0\mbox{ in }\mathbb{R}),$
also written as $B^{+}_{X}(x_{0},\varepsilon)$. (Of course, this local base can be made countable.) A finer $\mu$-metric $\mu^{\prime}\geqslant\mu$ (in 2.3(a)) gives a finer future topology, with more open sets.
@@
(37) $B-(x_{0},\varepsilon)=\{x\in X\,|\,\mu(x,x_{0})<\varepsilon\}\qquad\;\;(\varepsilon>0\mbox{ in }\mathbb{R}).$
However, we shall more often use the ‘reflective’ and ‘coreflective’ symmetric topologies, defined in 4.1 and 4.2, keeping a trace of the asymmetry of $\mu$ by a preorder, in 4.3.
(b) For a symmetric $\mu$-space $X$ the associated topological space ${\rm tp}(X)={\rm tp}^{+}(X)={\rm tp}^{-}(X)$ will generally be written $X$.
@@
(38) $\begin{array}[]{c}{\rm tp}^{+}\colon\mu_{\infty}\mathsf{Mtr}\to\mathsf{Top},\qquad{\rm tp}^{-}\colon\mu_{\infty}\mathsf{Mtr}\to\mathsf{Top},\\[5.0pt] {\rm tp}\colon!\mu_{\infty}\mathsf{Mtr}\to\mathsf{Top},\end{array}$
@@ 3. Basic examples and main properties
After examining $\mu$-structures on the line and recalling $\delta$-structures on the spheres, we deal with the main properties we are interested in: linear metrics, defined by a potential function (in 3.2), invariant metrics, defined by a relative norm (in 3.3), discrete and chaotic metrics (in 3.4), flat and symmetric metrics (in 3.5 and 3.6).
@@ 3.1. Metrics on the line and the spheres
(a) We shall generally test our notions on three basic structures, from which many others will be deduced later: the $\delta$-line $\delta\mathbb{R}$ (in (31)), the $\mu$-line $\mu\mathbb{R}$ (in 2.1(c)), and the Lawvere directed line $\delta_{0}\mathbb{R}$ (in (32)). Their future topology is determined by the following canonical local base at any point $x_{0}$:
- $\begin{array}[]{lll}\text{- for }\delta\mathbb{R}\colon&B^{+}(x_{0},\varepsilon)=[x_{0},x_{0}+\varepsilon[,\\[5.0pt] \text{- for }\mu\mathbb{R}\delta_{0}\mathbb{R}\colon&B^{+}(x_{0},\varepsilon)=\;]{\overline{\infty}},x_{0}+\varepsilon[\qquad&(\varepsilon>0).\end{array}$
This begins to make clear why we are using $\delta\mathbb{R}$ as the standard (metric) directed line: the ‘cone of the future’ at $x_{0}$ is appropriate to our goals, while that of $\delta_{0}\mathbb{R}$ is not. The $\mu$-line $\mu\mathbb{R}$ suits different goals, and its paths will be allowed to undergo losses and gains (as in the example of 0.1).
@@
(c) On the other hand, the orbit space $(\mu\mathbb{R})/\mathbb{Z}$ has the chaotic $\mu$-metric, always $\overline{\infty}$}, because two points $[x],[y]$ can be lifted to $\mu\mathbb{R}$ with arbitrarily small $\mu(x,y)$.
@@
Relative metric spaces lack a ‘standard $\mu$-circle’. As already said in the Introduction, we shall introduce a more general setting in Part III.
@@
As an ad hoc solution, one might introduce a notion of ‘local $\mu$-metric’, much in the same way as the chaotic preorder induced by the line on the circle can be replaced by a local preorder (see [G4], 1.9.3). Yet, this complicated framework would be ineffective for other drawbacks of relative metrics.
3.2. Linear $\mu$-metrics and potential functions
Let $(X,\mu)$ be a $\mu$-metric space.
(a) The $\mu$-metric is linear (see 2.2) if and only if it is finite and $\mu(x,y)=-\mu(y,x)$, for all pairs $x,y$. Moreover, $\mu(x,x)=0$, for all $x$.
@@ 3.2. Linear $\mu$-metrics and potential functions
Indeed, if $\mu$ is linear (which includes being finite, by definition), $\mu(x,y)+\mu(y,x)=\mu(x,x)=0$. Conversely, assuming that $\mu(x,y)=-\mu(y,x)$, the triangular inequality implies the reverse one.
@@
Let us note that the condition $\mu(x,y)=-\mu(y,x)$ does not imply that our $\mu$-metric is finite: for instance, on the set $\{0,1\}$ one can take $\mu(0,0)=0=\mu(1,1)$ and $\mu(0,1)=\infty=-\mu(1,0)$.
@@
(b) The $\mu$-metric is linear if and only if there is a function $\varphi\colon X\to\mathbb{R}$ (a potential of $\mu$) such that $\mu$ is its variation $\Delta\varphi$
(39) $\mu(x,y)=\Delta\varphi(x,y)=\varphi(y)-\varphi(x).$
@@
This expression plainly gives a linear $\mu$-metric. Conversely, if $\mu$ is linear, we define $\varphi(x)=\mu(x_{0},x)$, for a fixed point $x_{0}\in X$, and verify (39), using (a): $\varphi(y)-\varphi(x)=\mu(x_{0},y)-\mu(x_{0},x)=\mu(x_{0},y)+\mu(x,x_{0})=\mu(x,y).$
@@
Concretely, $\varphi$ could be the measure of any ‘variable’ on the ‘space’ $X$. Of course, if $\varphi$ is a $C^{1}$-function on a differentiable manifold, it is the potential of its gradient. One can view the variation $\mu=\Delta\varphi$ as a ‘finite-difference gradient’ of $\varphi$.
@@
The potential $\varphi$ is determined up to an additive constant (for $X\neq\varnothing$, of course): fixing an arbitrary point $x_{0}\in X$, we can always take $\varphi(x_{0})=0$ and
(40) $\varphi(x)=\mu(x_{0},x).$
@@
(c) Obvious examples can be deduced from the gravitational or electric fields, as in 0.1 and 6.6. A non-conservative field would produce a more general structure, studied in Part III.
@@
A product of linear $\mu$-spaces is not linear, generally: for instance $\mu\mathbb{R}^{2}$ is not. A linear $\delta$-metric is discrete: $\delta(x,y)=0$, for all $x,y$.
@@
(d) If $X$ and $Y$ are linear $\mu$-spaces, any $\mu$-map $f\colon X\to Y$ is an isometry. Any $\lambda$-Lipschitz map is an isometry $\lambda X\to Y$, and $\lambda$ is uniquely determined, unless $X$ has the chaotic $\delta$-metric, everywhere 0.
@@
In fact, from $\lambda\mu_{X}(x,x^{\prime})\geqslant\mu_{Y}(fx,fx^{\prime})$ and (a) we deduce that $\lambda\mu_{X}(x,x^{\prime})$ $\mu_{Y}(fx,fx^{\prime})$, so that $f$ is an isometry $\lambda X\to Y$. Since linear $\mu$-metrics are finite, $\lambda$ is uniquely determined unless $\mu_{X}$ is everywhere 0 (which includes the case $X=\varnothing$).
@@
(e) If $f\colon X\to Y$ is a $\lambda$-Lipschitz map from a linear $\mu$-space to a $\delta$-space, then:
@@
In fact, applying (a), the condition $\lambda\mu_{X}(x,x^{\prime})\geqslant\delta_{Y}(fx,fx^{\prime})\geqslant 0$ is equivalent to $\lambda\mu_{X}(x,x^{\prime})=0=\delta_{Y}(fx,fx^{\prime})$.
3.3. Invariance and relative norms
On a set $X$ acted on by a group $G$ we can consider a $\mu$-metric invariant up to the action of the group. More simply, we suppose that $X$ is an abelian group in additive notation, acting on itself.
(a) A $\mu$-metric $\mu$ on $X$ is said to be invariant up to translations if
(41) $\mu(x+z,y+z)=\mu(x,y)\qquad\qquad\qquad\qquad(x,y,z\in X).$
Here $\mu(x,x)=\mu(0,0)$, for all $x$: the metric is either regular or flat, according to the definitions of 2.2(b). In the second (less interesting) case, $\mu(x,y)=\pm\infty$ for all $x,y\in X$, as proved in 3.5.
(b) One can equivalently assign a relative norm, or $\mu$-norm, on $X$, which means a mapping ${\underline{\mu}}\colon X\to{\bf w}$ satisfying the axioms

### 13 May 2026 v2

@@ 2.5. Lipschitz maps
We say that a mapping $f\colon X\to Y$ between $\rho$-spaces is $\lambda$-Lipschitz, or a $\lambda$-map, if
(33) $\lambda\rho_{X}(x,x^{\prime})\geqslant\rho_{Y}(f(x),f(x^{\prime}))\qquad\qquad(x,x^{\prime}\in X),$
and we say that $f$ is a Lipschitz map, or $\rho_{\infty}$-map, if it is has some Lipschitz constant $\lambda\geqslant 0$, in the previous sense. Composing a $\lambda$-map with a consecutive $\lambda^{\prime}$-map we get a $\lambda\lambda^{\prime}$-map.
@@
The category $\rho\mathsf{Mtr}$ is a wide subcategory of the category $\rho_{\infty}\mathsf{Mtr}$ of $\rho$-spaces and Lipschitz maps. A Lipschitz isomorphism will be an isomorphism of the latter. The symmetric case is denoted as $\rho_{\infty}\mathsf{Mtr}$.
@@
(b) If $X$ is positive (a $\delta$-space) and $Y$ an arbitrary $\rho$-space, one can define, for every mapping $f\colon X\to Y$ (of sets) a Lipschitz weight, or seminorm $f||\in[0,\infty]$
(34) $||f||={\rm inf}\{\lambda\in[0,\infty[\,|\,\forall\;x,x^{\prime}\in X,\;\rho_{Y}(fx,fx^{\prime})\leqslant\lambda\delta_{X}(x,x^{\prime})\},$
@@
The categories $\delta_{\infty}\mathsf{Mtr}$ and $\delta_{\infty}\mathsf{Mtr}$ are multiplicatively weighted (or seminormed) by the Lipschitz weight. By this we mean that every morphism $f$ is assigned a weight $||f||\in[0,\infty[$ so that
@@
(c) The category $\rho_{\infty}\mathsf{Mtr}$ (resp. $\rho_{\infty}\mathsf{Mtr}$, $\delta_{\infty}\mathsf{Mtr}$, $\delta_{\infty}\mathsf{Mtr}$) is finitely complete and cocomplete, and the inclusion of $\rho\mathsf{Mtr}$ (resp. $\rho\mathsf{Mtr}$, $\delta\mathsf{Mtr}$, $\delta\mathsf{Mtr}$) preserves finite limits and colimits. But the ‘metric’ of a (co)limit in the larger category is only determined up to Lipschitz isomorphism.
If $X$ is a $\rho$-metric space and $\lambda\in[0,\infty[$, we shall write $\lambda X$ the same set equipped with the $\rho$-metric $\lambda\rho_{X}$. Thus, a $\lambda$-Lipschitz map is the same as a $\rho$-map $\lambda X\to Y$. This transformation preserves symmetric structures and positive structures.
@@ 2.6. Remarks
The extension to real valued metrics brings about aspects which strongly diverge from the positive case. We begin with some elementary observations. In particular, they show that the structure of $\rho\mathbb{R}$ highly restrains the Lipschitz maps defined on this space; in fact, we are more interested in maps from $\delta$-spaces to $\rho\mathbb{R}$ (as in 4.6).
(a) A $\rho_{\infty}$-map $f\colon\rho\mathbb{R}\to\rho\mathbb{R}$ is simply an affine map $f(x)=\lambda x+y_{0}$ (with $\lambda\geqslant 0$), and admits a unique Lipschitz constant, namely $\lambda$. (This will be extended to linear $\rho$-spaces, in 3.2(d).)
@@
(b) Any $\lambda$-Lipschitz maps $f\colon\rho\mathbb{R}\to\delta_{0}\mathbb{R}$ is constant, and $\lambda$ is necessarily 0 (see 3.2(e)).
@@
(c) Points as maps. A point of a $\rho$-space $X$ can be identified with a $\rho$-map $\delta\top\to X$ defined on the $\delta$-singleton. Formally, the forgetful functors $\rho_{\infty}\mathsf{Mtr}\to\mathsf{Set}$ and $\rho\mathsf{Mtr}\to\mathsf{Set}$ are represented by the $\delta$-singleton (the identity of the tensor product, see 6.1). If $X$ is positive every constant mapping $f\colon X\to Y$ is a $\rho$-map, since it can be factorised as $X\to\delta\top\to Y$.
On the other hand, a $\rho_{\infty}$-map $\rho\top\to Y$ can only reach a flat point of $Y$. (The $\rho$-singleton represents the ‘flat-part functor’ mentioned in 2.2(b).)
@@ 2.7. Future and past topologies
A $\rho$-space $X=(X,\rho)$ can be equipped with several topologies, which coincide in the symmetric case.
@@
(36) $B^{+}(x_{0},\varepsilon)=\{x\in X\,|\,\rho(x_{0},x)<\varepsilon\}\qquad\;\;(\varepsilon>0\mbox{ in }\mathbb{R}),$
also written as $B^{+}_{X}(x_{0},\varepsilon)$. (Of course, this local base can be made countable.) A finer $\rho$-metric $\rho^{\prime}\geqslant\rho$ (in 2.3(a)) gives a finer future topology, with more open sets.
@@
(37) $B-(x_{0},\varepsilon)=\{x\in X\,|\,\rho(x,x_{0})<\varepsilon\}\qquad\;\;(\varepsilon>0\mbox{ in }\mathbb{R}).$
However, we shall more often use the ‘reflective’ and ‘coreflective’ symmetric topologies, defined in 4.1 and 4.2, keeping a trace of the asymmetry of $\rho$ by a preorder (see 4.3).
(b) For a symmetric $\rho$-space $X$ the associated topological space ${\rm tp}(X)={\rm tp}^{+}(X)={\rm tp}^{-}(X)$ will generally be written $X$.
@@
(38) $\begin{array}[]{c}{\rm tp}^{+}\colon\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top},\qquad{\rm tp}^{-}\colon\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top},\\[5.0pt] {\rm tp}\colon!\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top},\end{array}$
@@ 3. Basic examples and main properties
After examining $\rho$-structures on the line and recalling $\delta$-structures on the spheres, we deal with the main properties we are interested in: linear metrics, defined by a potential function (in 3.2), and invariant metrics, defined by a real norm (in 3.3).
Several more technical properties of flat and symmetric $\rho$-spaces can be found in the preprint version of this article, Sections 3.5-3.6 [G5].
@@ 3.1. Metrics on the line and the spheres
(a) We shall generally test our notions on three basic structures, from which many others will be deduced later: the $\delta$-line $\delta\mathbb{R}$ (in (31)), the $\rho$-line $\rho\mathbb{R}$ (in 2.1(c)), and the Lawvere directed line $\delta_{0}\mathbb{R}$ (in (32)). Their future topology is determined by the following canonical local base at any point $x_{0}$:
- $\begin{array}[]{lll}\text{- for }\delta\mathbb{R}\colon&B^{+}(x_{0},\varepsilon)=[x_{0},x_{0}+\varepsilon[,\\[5.0pt] \text{- for }\rho\mathbb{R}\delta_{0}\mathbb{R}\colon&B^{+}(x_{0},\varepsilon)=\;]{\overline{\infty}},x_{0}+\varepsilon[\qquad&(\varepsilon>0).\end{array}$
This begins to make clear why we are using $\delta\mathbb{R}$ as the standard (metric) directed line: the ‘cone of the future’ at $x_{0}$ is appropriate to our goals, while that of $\delta_{0}\mathbb{R}$ is not. The $\rho$-line $\rho\mathbb{R}$ suits different goals, and its paths will be allowed to undergo losses and gains (as in the example of 0.2).
@@
(c) On the other hand, the orbit space $(\rho\mathbb{R})/\mathbb{Z}$ has the chaotic $\rho$-metric, always $\overline{\infty}$}, because two points $[x],[y]$ can be lifted to $\rho\mathbb{R}$ with arbitrarily small $\rho(x,y)$.
@@
Relative metric spaces lack a ‘standard $\rho$-circle’. As already said in the Introduction, we shall introduce a more general setting in Part III.
@@
As an ad hoc solution, one might introduce a notion of ‘local $\rho$-metric’, much in the same way as the chaotic preorder induced by the ordered line on the circle can be replaced by a local preorder (see [G4], 1.9.3). Yet, this complicated framework would be ineffective for other drawbacks of real metrics.
3.2. Linear real metrics and potential functions
Let $(X,\rho)$ be a $\rho$-metric space.
(a) The $\rho$-metric is linear (see 2.2) if and only if it is finite and $\rho(x,y)=-\rho(y,x)$, for all pairs $x,y$. Moreover, $\rho(x,x)=0$, for all $x$.
@@ 3.2. Linear real metrics and potential functions
Indeed, if $\rho$ is linear (which includes being finite, by definition), $\rho(x,y)+\rho(y,x)=\rho(x,x)=0$. Conversely, assuming that $\rho(x,y)=-\rho(y,x)$, the triangular inequality implies the reverse one.
@@
Let us note that the condition $\rho(x,y)=-\rho(y,x)$ does not imply that our $\rho$-metric is finite: for instance, on the set $\{0,1\}$ one can take $\rho(0,0)=0=\rho(1,1)$ and $\rho(0,1)=\infty=-\rho(1,0)$.
@@
(b) The $\rho$-metric is linear if and only if there is a function $\Phi\colon X\to\mathbb{R}$ (a potential of $\rho$) such that $\rho$ is its variation $\Delta\Phi$
(39) $\rho(x,y)=\Delta\Phi(x,y)=\Phi(y)-\Phi(x).$
@@
This expression plainly gives a linear $\rho$-metric. Conversely, if $\rho$ is linear, we define $\Phi(x)=\rho(x_{0},x)$, for a fixed point $x_{0}\in X$, and verify (39), using (a): $\Phi(y)-\Phi(x)=\rho(x_{0},y)-\rho(x_{0},x)=\rho(x_{0},y)+\rho(x,x_{0})=\rho(x,y).$
@@
Concretely, $\Phi$ could be the measure of any ‘variable’ on the ‘space’ $X$. Of course, if $\Phi$ is a $C^{1}$-function on a differentiable manifold, it is the potential of its differential 1-form. One can view the variation $\rho=\Delta\Phi$ as a ‘finite differential’ of $\Phi$.
@@
The potential $\Phi$ is determined up to an additive constant (for $X\neq\varnothing$, of course): fixing an arbitrary point $x_{0}\in X$, we can always take $\Phi(x_{0})=0$ and
(40) $\Phi(x)=\rho(x_{0},x).$
@@
(c) Obvious examples can be deduced from the gravitational or electric fields, as in 0.2 and 6.6. A non-conservative field would produce a more general structure, studied in Part III.
@@
A product of linear $\rho$-spaces is not linear, generally: for instance $\rho\mathbb{R}^{2}$ is not. A linear $\delta$-metric is discrete: $\delta(x,y)=0$, for all $x,y$.
@@
(d) If $X$ and $Y$ are linear $\rho$-spaces, any $\rho$-map $f\colon X\to Y$ is an isometry. Any $\lambda$-Lipschitz map is an isometry $\lambda X\to Y$, and $\lambda$ is uniquely determined, unless $X$ has the chaotic $\delta$-metric, everywhere 0.
@@
In fact, from $\lambda\rho_{X}(x,x^{\prime})\geqslant\rho_{Y}(fx,fx^{\prime})$ and (a) we deduce that $\lambda\rho_{X}(x,x^{\prime})$ $\rho_{Y}(fx,fx^{\prime})$, so that $f$ is an isometry $\lambda X\to Y$. Since linear $\rho$-metrics are finite, $\lambda$ is uniquely determined unless $\rho_{X}$ is everywhere 0 (which includes the case $X=\varnothing$).
@@
(e) If $f\colon X\to Y$ is a $\lambda$-Lipschitz map from a linear $\rho$-space to a $\delta$-space, then
@@
In fact, applying (a), the condition $\lambda\rho_{X}(x,x^{\prime})\geqslant\delta_{Y}(fx,fx^{\prime})\geqslant 0$ is equivalent to $\lambda\rho_{X}(x,x^{\prime})=0=\delta_{Y}(fx,fx^{\prime})$.
3.3. Invariance and real norms
On a set $X$ acted on by a group $G$ we can consider a $\rho$-metric invariant up to the action of the group. More simply, we suppose that $X$ is an abelian group in additive notation, acting on itself.
(a) A $\rho$-metric $\rho$ on $X$ is said to be invariant up to translations if
(41) $\rho(x+z,y+z)=\rho(x,y)\qquad\qquad\qquad\qquad(x,y,z\in X).$
Here $\rho(x,x)=\rho(0,0)$, for all $x$: the metric is either regular or flat, according to the definitions of 2.2(b) – also because $X$ is non-empty. In the second (less interesting) case, we have already seen that $\rho(x,y)=\pm\infty$, for all $x,y\in X$.
(b) One can equivalently assign a real norm, or $\rho$-norm, on $X$, which means a mapping ${\underline{\mu}}\colon X\to{\bf w}$ satisfying the axioms